OpenStudy (anonymous):

Sorry one more then im done, A merchant has coffee worth $20 a pound that she wished to mix with 70 pounds of coffee workth $80 a pound to get a mixture that can be sold for $40 a pound. How many pounds of the $20 coffee should be used??

OpenStudy (anonymous):

20x+80(70)=40(x+70) 20x+5600=40x+2800 40x-20x=5600-2800 20x=2800 x=140 pounds